# Automated Counting of Spanning Trees for Several Infinite families of Graphs

### v1

@@ 10 Conjectures and Further Discussion
Assuming that the previous conjecture holds, a natural question arises: given a graph family, is there a threshold for how many vertices you can "modify" (by edge deletion) while keeping the same B-Z constant? It turns out that Grid graphs and Torus Graphs have different B-Z constants, which may suggest that we can only modify a finite number of vertices. We state the resulting conjecture below and include definitions later.
@@
Let $\mathcal{G}:=\{G_{n}\}_{n}$ and :=$\mathcal{H}:=\{H_{n}\}_{n}$ be (indexed) graph families, such that $BZ(\mathcal{G})$ and $BZ(\mathcal{H})$ both exist. If the following hold:
@@
Let $\Delta(G)$ denote the maximum degree in a graph $G$. For two graphs $G,H$ such that $V(G)=V(H)$, we define $G\setminus H$ as the graph $V(G\setminus H):=V(G)$ and $E(G\setminus H):=E(G)\setminus E(H)$. Naturally, the conditions in the previous conjecture are satisfied when $\mathcal{G}=\mathcal{G}_{r}$ and $\mathcal{H}=\mathcal{H}_{r}$ (as defined in Section 2).

### v2

@@ 10 Conjectures and Further Discussion
Assuming that the previous conjecture holds, a natural question arises: given a graph family, is there a threshold for how many vertices you can "modify" (by edge deletion) while keeping the same B-Z constant? It turns out that Grid graphs and Torus Graphs have different B-Z constants, which may suggest that we can only modify a finite number of vertices. Originally, we considered the following conjecture. Note that the conditions in the following conjecture are satisfied when $\mathcal{G}=\mathcal{G}_{r}$ and $\mathcal{H}=\mathcal{H}_{r}$ (as defined in Section 2).
@@
(False) Let $\mathcal{G}:=\{G_{n}\}_{n}$ and :=$\mathcal{H}:=\{H_{n}\}_{n}$ be (indexed) families of connected graphs, such that $BZ(\mathcal{G})$ and $BZ(\mathcal{H})$ both exist. If the following hold:
@@
For two graphs $G,H$ such that $V(G)=V(H)$, we define $G\setminus H$ as the graph $V(G\setminus H):=V(G)$ and $E(G\setminus H):=E(G)\setminus E(H)$. Unfortunately, the conjecture above fails for families of sparse graphs, as proven by considering the following counterexample.
Counterexample. Let $\mathcal{H}$ be the family of graphs constructed in the proof of Proposition 4.0.2 with B-Z constant equal to $\frac{1}{2}$. In other words, $\mathcal{H}$ is the family of graphs resulting from subdividing each edge in a star graph once. From each $H_{n}\in\mathcal{H}$, construct $G_{n}$ by adding an edge between two non-leaf vertices. We see that $\tau(G_{n})=3\cdot\tau(H_{n})$, but $\mathcal{L}(G_{n})=\mathcal{L}(H_{n})$. Hence, $BZ(\mathcal{G})=\frac{1}{3}\cdot BZ(\mathcal{H})=\frac{1}{6}\neq\frac{1}{2}=BZ(\mathcal{H})$.